OpenStudy (anonymous):

f(x) = x + 1. Part 1: Find f(2). Part 2: Find x when f(x) = 2

OpenStudy (anonymous):

Part 1: Plug in 2 for x in the equation So, f(2)= 2+1 f(2)=3

OpenStudy (anonymous):

So, f = 1.5?

OpenStudy (anonymous):

Part 2: Set your equation f(x) equal to 2 and solve. So, x+1=2 subtract 1 from the x to isolate x. x=1

OpenStudy (anonymous):

lolwut?

OpenStudy (anonymous):

What do you not understand?

OpenStudy (anonymous):

So, f(2) =3 then that would come out to be f = 1.5?

OpenStudy (anonymous):

You are not solving for f.

OpenStudy (anonymous):

So thats the answer?

OpenStudy (anonymous):

f is the function, not a variable you want to find.

OpenStudy (anonymous):

@Guez, "f" is a function, not a multplier... for this problem, it's a rule that says, if you give me "x", I will give you back "x + 1"

OpenStudy (anonymous):

so if @Brittni0605 gives you x = 2, then your function gives back 2 + 1 = 3

OpenStudy (anonymous):

f(2)=3 and when f(x)= 2 your x value equals 1

OpenStudy (anonymous):

So the answers are.. Part 1 : f(2) = 3 Part 2 : x = 1 ? I understand how to do the first one but not the second one.

OpenStudy (anonymous):

Yes, correct

OpenStudy (anonymous):

Part 2 is like the reverse... you know your function gave back "2" as an answer... what must you have provided as an input to get "2" back?

OpenStudy (anonymous):

For the second part, the question is asking what x value will make 'x+1' = 2

OpenStudy (anonymous):

so lets say for part 2 it would have said Find x when f(x) = 4. The answer would have been x = 4?

OpenStudy (anonymous):

No, the answer would be 3

OpenStudy (anonymous):

If it said find f(x)=4: x+1=4 x=4-1 x=3

OpenStudy (anonymous):